OpenStudy (anonymous):

solve -2x-1=3x+5?2 for x

OpenStudy (diyadiya):

-2x-1=3x+ what?

OpenStudy (diyadiya):

5 or 2?

OpenStudy (anonymous):

oh no not that solve -2x-1=3x+5/2 for x

OpenStudy (diyadiya):

 x=-7/10